OpenStudy (anonymous):

Factor completely 10xy + 3y + 20ax + 6a.

OpenStudy (anonymous):

@iYuko

iYuko (iyuko):

10xy + 3y + 20ax + 6a = y(10x + 3) + 2a(10x + 3) = (y + 2a)(10x + 3)

iYuko (iyuko):

This site lags too much

iYuko (iyuko):

Medal Fan?